# Two-particle calculations with quantics tensor trains - solving the parquet equations

### v1

@@ IV.1 Quantics tensor train representation
Figure 2: Quantics representation and quantics tensor train of a univariate function
This representation can be generalized to functions of $N>1$ variables by applying the binary representation to each variable separately. For instance, a function $f(x,y,z)$ of three variables is represented as a tensor
@@
and the $L=NR$ indices are relabeled to $\sigma_{1}=x_{1},\sigma_{2}=y_{1},\sigma_{3}=z_{1},\sigma_{4}=x_{2},\ldots,\sigma_{L}=z_{R}$. Note that indices belonging to different variables are interleaved, which leads to an index ordering where all indices describing large length scales are grouped to the left, and all indices describing small length scales are grouped to the right. This tensor can then be factorized into a tensor train (TT), also known as matrix product state (MPS), of the form
@@ IV.2 Tensor Cross Interpolation (TCI)
TCI is more efficient than the SVD-based factorization, especially when the full tensor does not fit into the available memory [30, 34, 32]. SVD-based factorization requires reading all elements of the tensor, leading to an exponential growth of the computation time in $R$. In contrast, if the target tensor/function is low-rank, the computation time of the TCI-based factorization is linear in $R$, leading to an exponential speed-up over the SVD [34]. We refer the reader to Refs. 34, 32 for more technical details.
@@ IV.3 Matrix product operator (MPO)
As we will see in later sections, many operations in QTT can be implemented as the contraction of two MPOs illustrated in Fig. 3(b). The exact contraction will result in an MPS of large bond dimension $D_{\mathrm{A}}D_{\mathrm{B}}$, where $D_{\mathrm{A}}$ and $D_{\mathrm{B}}$ are the bond dimensions of the two input MPOs, respectively. Thus, the bond dimension of the resulting MPO must be truncated to some ${D_{\mathrm{max}}}$. The computational cost of a naive contraction followed by truncation scales $\mathcal{O}(D_{\mathrm{A}}^{3}D_{\mathrm{B}}^{3})$.
@@ V Parquet equations in QTT format
In order to evaluate the full set of parquet equations entirely within the QTT representation, we need to (i) represent the vertex functions in the QTT form; (ii) decompose Eqs. (12)-(17) into operations on QTTs. The latter can be implemented in a straightforward way by employing fundamental operations described in the previous section, namely MPO-MPO contractions. In the following subsections, we describe the quantics representation of vertex functions, check their compressibility to QTTs, and discuss the implementation of each step in solving the parquet equations. Two operations are particularly important:
@@ V.1 Quantics representation and compression of two-particle vertex functions
In this work, all functions represented in QTT format are functions of bosonic and fermionic frequencies, which are parameterized as $\nu=(2m-2^{R}+1)\pi/\beta$ and $\omega=(2m-2^{R})\pi/\beta$. The discrete index $m$ is then decomposed into quantics bits as in Eq. (19), and bits corresponding to different variables are then interleaved as illustrated in Fig. 4.
@@
The first step in using the QTT framework for solving the parquet equations is to investigate the compressibility of vertex functions in the above representation. We use the full vertex in the density channel ($F_{d}$) of the Hubbard atom for numerical demonstration. The fermionic frequency dependence of $F_{d}$ at $\omega=0$ is shown for various temperatures in Fig. 5. Note, that in the Hubbard atom the results are not separately dependent on temperature and $U$, but only on their ratio $\beta U$, since there are no other energy scales [35]. Before we focus on the scaling of bond dimension with temperature, let us first look at the bond dimension along the QTT at $\beta U=1$ for different grid sizes and tolerances set in TCI as shown in Fig. 6. Along the first and last bonds, the bond dimension grows exponentially as $D_{\ell}=\min(2^{\ell},2^{\mathcal{L}-\ell})$, which is the maximum bond dimension of an uncompressed factorization and represents maximum entanglement between these exponentially different length scales. In between, the bond dimension then saturates at a maximum bond dimension ${D_{\mathrm{max}}}$, indicating that the vertex structures are indeed compressible. The maximum bond dimension ${D_{\mathrm{max}}}$ is between $80$ and $400$, and increases with decreasing tolerance $\epsilon$. Since ${D_{\mathrm{max}}}$ is nearly independent of the grid parameter $R$, an exponential increase in the number of grid points ($\mathcal{O}(2^{R})$) can be achieved for linear cost ($\mathcal{O}(R)$) in runtime and memory. These findings are consistent with earlier results on the compressibility of vertices in Ref. 26.
@@
For large temperatures, such as $\beta U=1$ in the above example, the dominating structures are the diagonal and anti-diagonal part of the vertex. For small temperatures, i.e. large $\beta U$, the anti-diagonal vanishes and an additional cross structure appears for one of the Matsubara frequencies equal to ±$\pm\pi/\beta$. The source and physical meaning of different structures in two-particle vertices can be found in Ref. [1]. For large $\beta U$, the vertex is large even in some places where both Matsubara frequency arguments are large, as can be seen in Fig. 5(d). All in all, the structure of $F_{d}$ seems to be much simpler at large temperatures than at small ones, which leads us to expect an increase in bond dimension in QTT representation with increasing $\beta U$.
@@
In Fig. 7, we show ${D_{\mathrm{max}}}$ of the QTT representation of $F_{d}$ for various values of $\beta U$, different tolerances set in TCI, and various grid sizes. As expected from earlier considerations, the maximum bond dimension is quite low at high temperatures and steeply grows with $\beta U$ until $\beta U\approx 5.1$, where it reaches a maximum. Thereafter, the bond dimension decreases again, which does not conform with our expectations. The maximum can be related to the first global divergence of the irreducible vertex $\Gamma_{d}$ [35, 41, 42], which occurs at $\beta U=5.13715$ and is indicated by a dashed line in the plot. Although the irreducible vertex $\Gamma_{d}$ diverges at this value of $\beta U$, there is no phase transition connected with the divergence and $F_{d}$ remains finite. The presence of a maximum in ${D_{\mathrm{max}}}$ precisely at the first global divergence of $\Gamma_{d}$ is very interesting from two different perspectives. Firstly, we see that ${D_{\mathrm{max}}}$ stops growing with $\beta U$ and, thus, calculations for temperatures ranging from very low to very high are manageable within the QTT framework in this case. Secondly, although the full vertex $F_{d}$ does not contain singularities, we see fingerprints of this first global divergence of $\Gamma_{d}$ in the amount of length scale entanglement in the system represented by the maximum bond dimension. A deeper investigation of this behavior is left for future work.
@@
Figure 4: QTT representation for full vertex function.
Figure 6: Bond dimension $D_{l}$ at bond $l$ for different tolerances set in the TCI construction of the QTT of $F_{d}$ in the interleaved representation for $\beta=U=1$. The different values of $R$ correspond to different grid sizes ($2^{3R}$ grid points). The black line indicates the exponentially growing bond dimension of the full rank QTT without any truncation for $R=11$.
@@ V.2 Channel transformations
Figure 8: QTT implementation of the Bethe–Salpeter equations as tensor networks. (a) Channel transformation of a vertex in QTT form (blue) using an affine transform MPO (orange). This is described in more detail in Appendices B and C. (b) The Bethe–Salpeter equations are evaluated from QTT vertices are implemented using multiple MPO-MPO contractions (see text). The contraction itself is done in 4 steps: 1. Both QTT to be contracted are converted to MPOs. 2. The MPOs are contracted to a single MPO. 3. The duplicate $\omega_{\ell}^{\prime\prime}$ legs are removed. 4. The tensors with $\nu_{\ell}$ and $\nu^{\prime\prime}$ legs are factorized between their local legs, reaching the original QTT form.
@@
In this process, there are two distinct sources of error: the finite frequency box and the QTT truncation. As an example, Fig. 9(a) shows the error of a transformation of the full vertex $F$ in the ph channel parametrization to the pp channel parametrization (denoted as $F_{pp}$) in the $\nu,\nu^{\prime}$-plane. In two triangular regions, the upper right and lower left corner, errors are large due to the finite size of the frequency box, and are not caused by the QTT compression. These points correspond to grid points outside of the original frequency box that were transformed into the box by the ph to pp transformation. The missing data there is then replaced by zeros. This error can easily be reduced and shifted to higher Matsubara frequencies by increasing the size of the frequency box exponentially through increasing $R$. In the remaining diagonal region in between the two corners, the error is entirely due to the tensor train approximation. At a bond dimension of ${D_{\mathrm{max}}}=100$, the error is smaller than the error tolerance of $\epsilon=10^{-5}$ everywhere.
Figure 9: (a) Absolute normalized error of the ph to pp channel transformation of $F$ at $\omega=0$ in the fermionic Matsubara frequency plane for a $D_{max}=100$ and $R=7$, $\beta=U=1$, $\epsilon=10^{-8}$. Runtime of the ph to pp channel transformation of $F$ for various maximum bond dimensions and grid size parameters $R$ with $\beta=U=1$. The dashed lines indicate the cubic runtime increase with ${D_{\mathrm{max}}}$ (b) and linear increase by increasing $R$ (c), which corresponds to an exponential increase in the number of grid points.
@@ V.3 Parquet equation
The parquet equation (15) with its frequency shifts as in (16) can be solved entirely in QTT using the methods introduced in the previous section to convert all the vertex functions $\Lambda,\Phi$ to the required channel and performing their summation and subtraction as described in Ref. [32]. The resultant QTTs for $F$ (and subsequently $\Gamma$ obtained from (13)) are then compressed to a maximum bond dimension ${D_{\mathrm{max}}}$ in $\mathcal{O}({D_{\mathrm{max}}}^{3}R)$ computation time. Hence, the parquet equation has the same $\mathcal{O}({D_{\mathrm{max}}}^{3}R)$ computational cost as channel transformations. Further investigation of error and runtime scaling can be found in App. D.
@@ V.4 Bethe–Salpeter equation
The costliest part of the iterative parquet scheme are the Bethe–Salpeter equations (12), where two infinite Matsubara sums have to be performed. These can be implemented as a sequence of matrix multiplications as $(\Gamma\chi_{0})F$. At each multiplication step, we have to perform the product of two vertex functions $A$ and $B$ as
@@
We recast this operation into a matrix multiplication as
$C^{\nu\nu^{\prime\prime}\omega}$ $=\sum_{\nu^{\prime}\omega^{\prime}}\tilde{A}^{\nu\nu^{\prime}\omega\omega^{\prime}}\tilde{B}^{\nu^{\prime}\nu^{\prime\prime}\omega^{\prime}\omega^{\prime\prime}}\Big{|}_{\omega=\omega^{\prime\prime}},$ (24)
$\tilde{A}^{\nu\omega}_{\nu^{\prime}\omega^{\prime}}$ $:=A^{\nu\nu^{\prime}\omega}\delta_{\omega,\omega^{\prime}},$ (25)
$\tilde{B}^{\nu^{\prime}\omega^{\prime}}_{\nu^{\prime\prime}\omega^{\prime\prime}}$ $:=B^{\nu^{\prime}\nu^{\prime\prime}\omega^{\prime}}\delta_{\omega^{\prime},\omega^{\prime\prime}},$ (26)
where $|_{\omega=\omega^{\prime\prime}}$ denotes the restriction of the result to $\omega=\omega^{\prime\prime}$. Figure 8(b) illustrates the implementation of these equations in QTT. Note that dummy indices $\omega^{\prime}$ have to be introduced, such that the $\omega$ tensors are diagonal in their two local legs. We also treat $\chi_{0}$ as a vertex object with two fermionic frequency axes and one bosonic frequency axis.
@@
After each MPO-MPO contraction, the bond dimensions are truncated to ${D_{\mathrm{max}}}$. If all MPOs are trucated to ${D_{\mathrm{max}}}$, the computational cost of each contraction is expected to scale as $\mathcal{O}({D_{\mathrm{max}}}^{4}R)$ as described in Sec. IV.3, which is more expensive than the channel transformation for large ${D_{\mathrm{max}}}$.
@@
We now conduct numerical tests to verify the accuracy of the operation and the scaling of the computational cost. Figure 10(a) shows the maximum absolute normalized error [cf. Eq. (22)] of the QTT implementation of the BSE in the density channel for various grid sizes depending on ${D_{\mathrm{max}}}$. The dashed lines indicate the results applying these matrix multiplications for the full numerical data and without the compressed QTTs. The error of these “dense grid calculations” is due to the finite size of the grid and, thus, caused by the truncated Matsubara sum. The results can be improved by increasing the grid size.
@@
For the dense grid calculations, the improved results come at high cost since increasing the grid parameter $R$ leads to an exponential increase in memory and computational cost. On the contrary, when using QTTs the memory and computational cost only increase linearly with $R$, which can be observed in Fig. 10(c), where the linear dependence of the runtime of the BSE on $R$ is shown for different maximum bond dimensions.
@@
Moreover, it can be observed that the QTT BSE errors converge to the box results. Interestingly, for larger grid sizes, slightly larger ${D_{\mathrm{max}}}$ are needed to reach the same error level as in the case of smaller boxes. However, a maximum normalized error of $\mathcal{O}(10^{-4})$ can be easily reached using QTTs for a still reasonable bond dimension of 200. Without the use of QTTs this would correspond to calculations with objects of $8\times 2^{3\times 12}\simeq 5.5\times 10^{11}$ bytes, for which multiple nodes on a cluster would need to be occupied. With the current SVD-based QTT matrix multiplication implementation the BSE operation only takes about 400 seconds on a single 512 GB node without parallelization.
@@
Figure 10: (a) Maximum absolute normalized error $||\Delta x||_{\infty}/||x||_{\infty}$ of the BSE with QTCI for the reducible vertex $\Phi_{d}$ compared to the exact values for various grid sizes depending on the maximum bond dimension with $U=\beta=1$ and $\epsilon=10^{-10}$. The dashed lines indicate the results from the dense grid calculation without QTCI. Runtime of the BSE for $\Phi_{d}$ for various maximum bond dimensions and grid size parameters $R$ with $\beta=U=1$. The dashed lines indicate the quartic runtime increase with ${D_{\mathrm{max}}}$ (b) and linear increase by increasing $R$ (c), which corresponds to an exponential increase in the number of grid points.
@@
V.5 Schwinger-Dyson equation
Similarly to the BSE evaluation, two infinite Matsubara sums have to be performed in the SDE in Eq. (17). Hence, qualitatively similar results to the BSE case are expected. This is indeed the case. The results are presented in Fig. 15 in Appendix E.

### v2

@@ IV.1 Quantics tensor train representation
Figure 2: Quantics representation and quantics tensor train of a univariate function.
This representation can be generalized to functions of $N>1$ variables by applying the binary representation to each variable separately. For instance, a function $f(x,y,z)$ of three variables is represented as a tensor depending on $3R$ binary indices
@@
and the $L=3R$ indices are relabeled to $\sigma_{1}=x_{1},\sigma_{2}=y_{1},\sigma_{3}=z_{1},\sigma_{4}=x_{2},\ldots,\sigma_{L}=z_{R}$. Note that indices belonging to different variables are interleaved, which leads to an index ordering where all indices describing large length scales are grouped to the left, and all indices describing small length scales are grouped to the right. This tensor can then be factorized into a tensor train (TT), also known as matrix product state (MPS), of the form
@@ IV.2 Tensor Cross Interpolation (TCI)
TCI is more efficient than the SVD-based factorization, especially when the full tensor does not fit into the available memory [32, 36, 34]. SVD-based factorization requires reading all elements of the tensor, leading to an exponential growth of the computation time in $R$. In contrast, if the target tensor/function is low-rank, the computation time of the TCI-based factorization is linear in $R$, leading to an exponential speed-up over the SVD [36]. We refer the reader to Refs. 36, 34 for more technical details.
@@ IV.3 Matrix product operator (MPO)
As we will see in later sections, many operations in QTT can be implemented as the contraction of two MPOs, illustrated in Fig. 3(b). The exact contraction will result in an MPS of large bond dimension $D_{\mathrm{A}}D_{\mathrm{B}}$, where $D_{\mathrm{A}}$ and $D_{\mathrm{B}}$ are the bond dimensions of the two input MPOs, respectively. Thus, the bond dimension of the resulting MPO must be truncated to some $D_{\mathrm{max}}$. The computational cost of a naive contraction followed by truncation scales $\mathcal{O}(D_{\mathrm{A}}^{3}D_{\mathrm{B}}^{3})$.
@@ V Parquet equations in QTT format
In order to evaluate the full set of parquet equations completely within the QTT representation, we need to (i) represent the vertex functions in the QTT form; and (ii) decompose Eqs. (12)-(17) into operations on QTTs. The latter can be implemented in a straightforward way by employing fundamental operations described in the previous section, namely MPO-MPO contractions. In the following subsections, we describe the quantics representation of vertex functions, check their compressibility to QTTs, and discuss the implementation of each step in solving the parquet equations. Two operations are particularly important:
@@ V.1 Quantics representation and compression of two-particle vertex functions
In this work, all functions represented in QTT format are functions of bosonic and fermionic frequencies, which are parameterized as $\nu=(2m-2^{R}+1)\pi/\beta$ and $\omega=(2m-2^{R})\pi/\beta$, respectively. The discrete index $m\in\{0,\ldots,2^{R}-1\}$ is then decomposed into quantics bits as in Eq. (19), and bits corresponding to different variables are then interleaved as illustrated in Fig. 4.
Figure 4: QTT representation for full vertex function.
@@
The first step in using the QTT framework for solving the parquet equations is to investigate the compressibility of vertex functions in the above representation. We use the full vertex in the density channel ($F_{d}$) of the Hubbard atom for numerical demonstration. The fermionic frequency dependence of $F_{d}$ at $\omega=0$ is shown for various temperatures in Fig. 5. Note, that in the Hubbard atom the results are not separately dependent on temperature and $U$, but only on their ratio $\beta U$, since there are no other energy scales [37]. Before we focus on the scaling of bond dimension with temperature, let us first look at the bond dimension along the QTT at $\beta U=1$ for different grid sizes and tolerances set in TCI as shown in Fig. 6. Moving inward from the first and last bonds, the bond dimension grows exponentially as $D_{\ell}=\min(2^{\ell},2^{\mathcal{L}-\ell})$, which is the maximum bond dimension of an uncompressed factorization and represents maximum entanglement between these exponentially different length scales. In between, the bond dimension then saturates at a maximum bond dimension $D_{\mathrm{max}}$, therefore indicating that the vertex structures are indeed compressible. The maximum bond dimension $D_{\mathrm{max}}$ is between $80$ and $400$, and increases with decreasing tolerance $\epsilon$. Importantly, $D_{\mathrm{max}}$ is nearly independent of the grid parameter $R$, an exponential increase in the number of grid points ($\mathcal{O}(2^{R})$) can be achieved for linear cost ($\mathcal{O}(R)$) in runtime and memory. These findings are consistent with earlier results on the compressibility of vertices in Ref. 28.
@@
For large temperatures, such as $\beta U=1$ in the above example, the dominating structures are the diagonal and anti-diagonal part of the vertex. For small temperatures, i.e. large $\beta U$, the anti-diagonal vanishes and an additional cross structure appears for one of the Matsubara frequencies equal to ±$\pm\pi/\beta$. The source and physical meaning of different structures in two-particle vertices have been discussed in Ref. [1]. For large $\beta U$, the vertex is large even in some places where both Matsubara frequency arguments are large, as can be seen in Fig. 5(d). All in all, the structure of $F_{d}$ seems to be much simpler at large temperatures than at small ones, which leads us to expect an increase in bond dimension in QTT representation with increasing $\beta U$.
Figure 6: Bond dimension $D_{\ell}$ at bond $\ell$ for different tolerances set in the TCI construction of the QTT of $F_{d}$ in the interleaved representation for $\beta=U=1$. The different values of $R$ correspond to different grid sizes ($2^{3R}$ grid points). The black line indicates the exponentially growing bond dimension of the full rank QTT without any truncation for $R=11$.
@@
In Fig. 7, we show $D_{\mathrm{max}}$ of the QTT representation of $F_{d}$ for various values of $\beta U$, different tolerances set in TCI, and various grid sizes. As expected from earlier considerations, the maximum bond dimension is quite low at high temperatures and steeply grows with $\beta U$ until $\beta U\approx 5.1$, where it reaches a maximum. Thereafter, the bond dimension decreases again, which does not conform with our expectations. The maximum can be related to the first global divergence of the irreducible vertex $\Gamma_{d}$ [37, 43, 44, 45, 46], which occurs at $\beta U=5.13715$ and is indicated by a dashed line in the plot. Although the irreducible vertex $\Gamma_{d}$ diverges at this value of $\beta U$, there is no phase transition connected with the divergence and $F_{d}$ remains finite. The presence of a maximum in $D_{\mathrm{max}}$ precisely at the first global divergence of $\Gamma_{d}$ is very interesting from two different perspectives. Firstly, we see that $D_{\mathrm{max}}$ stops growing with $\beta U$ and, thus, calculations for temperatures ranging from very low to very high are manageable within the QTT framework in this case. Secondly, although the full vertex $F_{d}$ does not contain singularities, we see fingerprints of this first global divergence of $\Gamma_{d}$ in the amount of length scale entanglement in the system represented by the maximum bond dimension. A deeper investigation of this behavior is left for future work.
@@ V.2 Channel transformations
Figure 8: QTT implementation of the Bethe–Salpeter equations as tensor networks. (a) Channel transformation of a vertex in QTT form (blue) using an affine transform MPO (orange). This is described in more detail in Appendices B and C. (b) The Bethe–Salpeter equations are evaluated from QTT vertices and are implemented using multiple MPO-MPO contractions (see text). The contraction itself is done in 4 steps: 1. Both QTT to be contracted are converted to MPOs. 2. The MPOs are contracted to a single MPO. 3. The duplicate $\omega_{\ell}^{\prime\prime}$ legs are removed. 4. The tensors with $\nu_{\ell}$ and $\nu^{\prime\prime}$ legs are factorized between their local legs, reaching the original QTT form.
@@
In this process, there are two distinct sources of error: the finite frequency box and the QTT truncation. As an example, Fig. 9(a) shows the absolute normalized error ($||\Delta F_{\textit{pp}}||:=|F_{\textit{pp},\mathrm{trafo}}-F_{\textit{pp},\mathrm{exact}}|/||F_{\textit{pp},\mathrm{exact}}||_{\infty}$) of a transformation of the full vertex $F$ in the ph channel parametrization to the pp channel parametrization (denoted as $F_{\textit{pp}}$) in the $\nu,\nu^{\prime}$-plane. In two triangular regions, the upper right and lower left corner, errors are large due to the finite size of the frequency box, and are not caused by the QTT compression. These points correspond to grid points outside of the original frequency box that were transformed into the box by the ph to pp transformation. The missing data there can be either replaced by zeros (which corresponds to open boundary conditions of the affine transformation, see App. B) or by values extrapolated from another part of the frequency box (e.g. by using periodic boundary conditions for the affine transformation, as in Ref. [47, footnote 14]). We checked that for the examples shown in this paper the average difference in error between the two options is small. We used periodic boundary conditions for all results presented in the manuscript. For the half-filled Hubbard atom this leads to better representation of missing values on the diagonal due to high symmetry of the full vertex in this case. In general, the choice of boundary conditions can be adapted to the problem at hand. Since the error can easily be reduced and shifted to higher Matsubara frequencies by increasing the size of the frequency box exponentially through increasing $R$, we do not expect the choice of boundary conditions to have significant effect on the error.
In the remaining diagonal region in between the two yellow corners in Fig. 9(a), the error is entirely due to the tensor train approximation. At a bond dimension of $D_{\mathrm{max}}=100$, the error is smaller than the error tolerance of $\epsilon=10^{-5}$ everywhere.
Figure 9: (a) Absolute normalized error $||\Delta F_{\textit{pp}}||:=|F_{\textit{pp},\mathrm{trafo}}-F_{\textit{pp},\mathrm{exact}}|/||F_{\textit{pp},\mathrm{exact}}||_{\infty}$ of the ph to pp channel transformation of $F$ at $\omega=0$ in the fermionic Matsubara frequency plane for a $D_{max}=100$ and $R=7$, $\beta=U=1$, $\epsilon=10^{-8}$. (b)–(c) Runtime of the ph to pp channel transformation of $F$ for various maximum bond dimensions and grid size parameters $R$ with $\beta=U=1$. The dashed lines indicate (b) the cubic runtime increase with $D_{\mathrm{max}}$, and (c) the linear increase by increasing $R$, which corresponds to an exponential increase in the number of grid points.
@@ V.3 Parquet equation
The parquet equation (15) with its frequency shifts as in (16) can be solved entirely in QTT by first converting all the vertex functions $\Lambda,\Phi$ to the required channel as described in the previous section, then performing their summation and subtraction as shown in Ref. [34, Sec. 4.7]. The resultant QTTs for $F$ (and subsequently $\Gamma$ obtained from (13)) are then compressed to a maximum bond dimension $D_{\mathrm{max}}$ in $\mathcal{O}(D_{\mathrm{max}}^{3}R)$ computation time. Hence, the parquet equation has the same $\mathcal{O}(D_{\mathrm{max}}^{3}R)$ computational cost as channel transformations. Further investigation of error and runtime scaling can be found in App. D.
@@ V.4 Bethe–Salpeter equation
The costliest part of the iterative parquet scheme are the Bethe–Salpeter equations (12), where two infinite Matsubara sums have to be performed. These can be implemented as a sequence of matrix multiplications as $(\Gamma\chi_{0})F$, where $\chi_{0}$ is treated as a vertex object with two fermionic frequency axes and one bosonic frequency axis. At each multiplication step, we have to compute the product of two vertex functions $A$ and $B$ as
@@
To express this summation as matrix multiplication, we introduce dummy indices $\omega^{\prime}$ and $\omega^{\prime\prime}$, such that
$C^{\nu\nu^{\prime\prime}\omega}$ $=\sum_{\nu^{\prime}\omega^{\prime}}\tilde{A}^{\nu\omega}_{\nu^{\prime}\omega^{\prime}}\tilde{B}^{\nu^{\prime}\omega^{\prime}}_{\nu^{\prime\prime}\omega^{\prime\prime}}\Big{|}_{\omega=\omega^{\prime\prime}},$ (24a)
$\tilde{A}^{\nu\omega}_{\nu^{\prime}\omega^{\prime}}$ $:=A^{\nu\nu^{\prime}\omega}\delta_{\omega,\omega^{\prime}},$ (24b)
$\tilde{B}^{\nu^{\prime}\omega^{\prime}}_{\nu^{\prime\prime}\omega^{\prime\prime}}$ $:=B^{\nu^{\prime}\nu^{\prime\prime}\omega^{\prime}}\delta_{\omega^{\prime},\omega^{\prime\prime}},$ (24c)
where $|_{\omega=\omega^{\prime\prime}}$ denotes the restriction of the result to $\omega=\omega^{\prime\prime}$. Note that Eq. (24a) has the structure of a matrix multiplication in the combined index $(\nu^{\prime},\omega^{\prime})$. Thus, this equation can be evaluated in QTT format through standard MPO-MPO contraction, as is illustrated in Fig. 8(b). Introducing dummy indices has a runtime and memory cost of $\mathcal{O}(D_{\mathrm{max}}^{2}R)$, which is much smaller than the cost of other steps in the algorithm.
@@
After each MPO-MPO contraction, the bond dimensions are truncated to $D_{\mathrm{max}}$. If all MPOs are truncated to $D_{\mathrm{max}}$, the computational cost of each contraction is expected to scale as $\mathcal{O}(D_{\mathrm{max}}^{4}R)$ as described in Sec. IV.3, which is more expensive than the channel transformation for large $D_{\mathrm{max}}$.
@@
We now conduct numerical tests to verify the accuracy of the operation and the scaling of the computational cost. Figure 10(a) shows the dependence on $D_{\mathrm{max}}$ of the maximum absolute normalized error ($||\Delta\Phi_{d}||_{\infty}:=||\Phi_{d,\mathrm{BSE}}-\Phi_{d,\mathrm{exact}}||_{\infty}/||\Phi_{d,\mathrm{exact}}||_{\infty}$) [cf. Eq. (22)] of the QTT implementation of the BSE in the density channel for various grid sizes. The dashed lines indicate the results applying these matrix multiplications for the full numerical data and without the compressed QTTs. The error of these “dense grid calculations” is due to the finite size of the grid and, thus, caused by the truncated Matsubara sum. The results can be improved by increasing the grid size.
@@
For the dense grid calculations, the improved results come at high cost since increasing the grid parameter $R$ leads to an exponential increase in memory and computational cost. By contrast, when using QTTs the memory and computational cost only increase linearly with $R$, which can be observed in Fig. 10(c), where the linear dependence of the runtime of the BSE on $R$ is shown for different maximum bond dimensions.
@@
Moreover, it can be observed that the QTT BSE errors converge to the box results. Interestingly, for larger grid sizes, slightly larger $D_{\mathrm{max}}$ are needed to reach the same error level as in the case of smaller boxes. However, a maximum normalized error $<10^{-3}$ can be easily reached using QTTs for a still reasonable bond dimension of 200. Without the use of QTTs this would correspond to calculations with objects of $8\times 2^{3\times 12}\simeq 5.5\times 10^{11}$ bytes, for which multiple nodes on a cluster would need to be occupied. With the current SVD-based QTT matrix multiplication implementation the BSE operation only takes about 400 seconds on a single 512 GB node (equipped with two AMD EPYC 7713 processors) without parallelization.
@@
Figure 10: (a) Dependence on $D_{\mathrm{max}}$ of the maximum absolute normalized error $||\Delta\Phi_{d}||_{\infty}:=||\Phi_{d,\mathrm{BSE}}-\Phi_{d,\mathrm{exact}}||_{\infty}/||\Phi_{d,\mathrm{exact}}||_{\infty}$ of the BSE with QTCI for the reducible vertex $\Phi_{d}$ compared to the exact values, for various grid sizes, with $U=\beta=1$ and $\epsilon=10^{-10}$. The dashed lines indicate the results from the dense grid calculation without QTCI. (b)–(c) Runtime of a single evaluation of the BSE for $\Phi_{d}$ for various maximum bond dimensions and grid size parameters $R$ with $\beta=U=1$. The dashed lines indicate (b) the quartic runtime increase with $D_{\mathrm{max}}$, and (c) the linear increase by increasing $R$, which corresponds to an exponential increase in the number of grid points.
@@
V.5 Schwinger–Dyson equation
Similarly to the BSE evaluation, two infinite Matsubara sums have to be performed in the SDE in Eq. (17). Hence, qualitatively similar scaling with $D_{\mathrm{max}}$ and $R$ to the BSE case are expected. This is indeed the case. The results are presented in Fig. 16 in Appendix E.